### v2

@@ 3 A Generic Approach for Beyond the Worst-Case Analysis
Fix a metric space $(X,\delta)$ and an algorithm ${\mathcal{A}}$ for online metric matching. Then, there exists an algorithm ${\mathcal{A}}^{\prime}$ such that for all set of points $I$ given before the arrival of any request, set of servers $S$, and set of requests $R$ with $|S|\geq|I|=|R|$, $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)$. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
@@
We construct the algorithm ${\mathcal{A}}^{\prime}$ as follows. Fix three sets of points $S,I,R$ as required. Let $M^{S,I}$ be the matching between $S$ and $I$ with the minimum cost that matches all points in $I$. Then, we run ${\mathcal{A}}$ with the set of initial servers being $I$. For each new request $r$, suppose that ${\mathcal{A}}$ matches it to $i\in I$, and ${\mathcal{A}}^{\prime}$ will match $r$ to $M^{S,I}(i)\in S$.
@@
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\sum_{i\in I}\delta(M^{S,I}(i),i)+\mathrm{cost}_{{\mathcal{A}}}(I,R)=\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\,,$
@@
3.1 Stochastic Setting
We first apply Theorem 1 to the case where all request locations are independently drawn from a known distribution $\mathbb{D}$. We establish in the following theorem that an algorithm for stochastic servers and requests can be transformed into an algorithm for adversarial servers and stochastic requests in a black-box manner. In particular, for balanced markets, this transformation preserves the competitive ratio of the algorithm up to a constant factor.
@@ 3.1 Stochastic Setting
Given a metric space $(X,\delta)$ and a distribution $\mathbb{D}$ over $X$, assume the existence of an algorithm ${\mathcal{A}}$ such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$. Then, an algorithm ${\mathcal{A}}^{\prime}$ exists such that $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $|S|\geq n$. If we further have $|S|=n$, then $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq(2\alpha+1)\cdot\mathrm{OPT}(S,n)$.
The rest of this subsection is devoted to proving Theorem 2. We start with proving the triangle inequality for $\mathrm{OPT}(\cdot,\cdot)$.
@@
Given $S,I,R$ with $|S|=|I|=|R|$, let $M_{1}$ be the min-cost perfect matching between $S$ and $I$, and let $M_{2}$ be the min-cost perfect matching between $I$ and $R$. Note that
@@
Assuming that requests are stochastic, we start by lower bounding the optimal offline cost in balanced markets with adversarial servers. In particular, we show that it is at least the optimal offline cost in balanced markets with stochastic servers and requests up to a constant factor.
@@
For all metric space $(X,\delta)$, distribution $\mathbb{D}$ over $X$, and set of servers $S$ with $|S|=n$, $\mathrm{OPT}(S,n)\geq\frac{1}{2}\cdot\mathrm{OPT}(n)$.
@@
Given a set of servers $S$ with $|S|=n$, by Lemma 6, we know that $\mathrm{OPT}(I,R)\leq\mathrm{OPT}(I,S)+\mathrm{OPT}(S,R)$ for all sets of points $I,R$ with $|I|=|R|=n$. Taking expectations on both sides over $I,R\sim\mathbb{D}^{n}$, we obtain
$\mathrm{OPT}(n)\leq\mathbb{E}_{I,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(I,S)+\mathrm{OPT}(S,R)\right]=2\cdot\mathrm{OPT}(S,n)\,,$
@@
Now, we are ready to prove Theorem 2. Given a metric space $(X,\delta)$ and a distribution $\mathbb{D}$ over $X$, let ${\mathcal{A}}$ be an algorithm such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$, and we construct an algorithm ${\mathcal{A}}^{\prime}$ for adversarial servers and stochastic requests as follows. Given as input a set of servers $S$ and a set of requests $R$ with $|S|\geq|R|=n$, we randomly generate a set of points $I\sim\mathbb{D}^{n}$ before the arrival of any request, and let ${\mathcal{A}}^{\prime}$ be the algorithm as per Theorem 2, which follows that
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\,.$
@@
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathbb{E}_{I,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\right]=\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)\,,$
which concludes the first part in Theorem 2. Furthermore, when $|S|=n$, by the guarantee of ${\mathcal{A}}$ and Lemma 7, we obtain
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\alpha\cdot\mathrm{OPT}(n)\leq(2\alpha+1)\cdot\mathrm{OPT}(S,n)\,,$
concluding the second part in Theorem 2.
3.2 Prediction Setting
In this subsection, we focus on balanced markets, i.e., $m=n$, and assume that the locations of servers and requests are adversarially chosen. In addition, we assume that algorithms are equipped with a prediction of request locations. We then apply Theorem 1 to give an algorithm whose performance smoothly depends on the prediction error, i.e., a consistent algorithm. More formally, the prediction is given in the form of a set of predicted request locations $R^{\prime}$ with $|R^{\prime}|=n$, which is provided before the arrival of any request, and the prediction error is defined as $\eta\triangleq\mathrm{OPT}(R^{\prime},R)$. We reiterate that the prediction error is not known by algorithms. Throughout, we slightly abuse notations and use $\mathrm{OPT}\triangleq\mathrm{OPT}(S,R)$ to denote the optimal offline cost.
@@ 3.2 Prediction Setting
Given a metric space $(X,\delta)$, assume that there exists an $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions. Then, there exists an algorithm ${\mathcal{A}}^{\prime}$ such that, when equipped with a prediction of request locations with error $\eta$, ${\mathcal{A}}^{\prime}$ is $(1+\frac{\eta(\alpha+1)}{\mathrm{OPT}})$-competitive. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
@@
We construct ${\mathcal{A}}^{\prime}$ as follows. Given as input a set of servers $S$, a set of requests $R$, and a prediction of request locations $R^{\prime}$ with $|S|=|R|=|R^{\prime}|$ and $\eta=\mathrm{OPT}(R,R^{\prime})$, let ${\mathcal{A}}^{\prime}$ be the algorithm as per Theorem 2 by setting $I=R^{\prime}$. It gives that,
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)$ $\leq\mathrm{OPT}(S,R^{\prime})+\mathrm{cost}_{{\mathcal{A}}}(R^{\prime},R)\leq\mathrm{OPT}(S,R)+\mathrm{OPT}(R,R^{\prime})+\mathrm{cost}_{{\mathcal{A}}}(R^{\prime},R)$
4 Stochastic Setting: Euclidean Space with Smooth Distributions
In this section, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance and $\mathbb{D}$ to be $\beta$-smooth [^4] The smoothness of distributions is a mild technical assumption allowing us to lower bound $\mathrm{OPT}(n)$. To see the necessity, consider $\mathbb{D}$ a non-smooth distribution supported on a singleton, in which case $\mathrm{OPT}(n)=0$. for some constant $\beta>0$, and we will apply Theorem 2 to derive algorithms for both balanced and unbalanced markets. In Section 4.1, we focus on uniform distribution, which is a special case of smooth distributions, and give an $O(1)$-competitive algorithm for all $d\geq 1$ in balanced markets.
Next, we consider balanced markets with smooth distributions in Section 4.2. Recall that when $d=1$, an $O(1)$-competitive algorithm in balanced markets for general distributions is given by [GGPW19]. Moreover, this algorithm is $O((\log\log\log n)^{2})$-competitive for general metrics and distributions. However, for $d=2$ and smooth distributions, we are only able to estimate the optimal offline cost up to a $\Theta(\sqrt{\log n})$ factor, which prevents our approach from achieving an $o(\sqrt{\log n})$ competitive ratio. Hence, we only focus on the case with $d\geq 3$, where we present an $O(1)$-competitive algorithm.
Finally, in Section 4.3, we extend our results for smooth distributions to unbalanced markets that possess at most linearly many excess servers, i.e., $n\leq m\leq O(n)$. We give an algorithm with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
4.1 Prelude: Balanced Markets with Uniform Distribution
In this subsection, we consider balanced markets and assume $\mathbb{D}$ to be uniform distribution. In the following theorem, we show that combining Theorem 2 with the Hierarchical Greedy algorithm of [Kan22] results in an $O(1)$-competitive algorithm for all $d\geq 1$.
Theorem 4.
Suppose that $m=n$. For $d\geq 1$, an $O(1)$-competitive algorithm exists for adversarial servers and uniformly distributed requests.
Proof.
Recall that [Kan22] gives an $O(1)$-competitive algorithm for $d\geq 1$ when both servers and requests follow uniform distribution. Combining the above algorithm with Theorem 2 concludes the proof. ∎
@@ 4.2 Balanced Markets
In this subsection, we assume $d\geq 3$ and consider balanced markets with smooth distributions. We will prove that the SOAR algorithm is $O(1)$-competitive for $d\geq 3$ when servers and requests are randomly drawn from $\mathbb{D}$. Together with Theorem 2, this leads to an $O(1)$-competitive algorithm for $d\geq 3$ with adversarial servers and stochastic requests.
@@
Suppose that $m=n$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$, an $O(1)$-competitive algorithm exists for $d\geq 3$.
The rest of this subsection is devoted to proving Theorem 5. By Theorem 2, it suffices to show that the SOAR algorithm is $O(1)$-competitive for $d\geq 3$ when both servers and requests are randomly drawn from $\mathbb{D}$. We compute the expected cost of the SOAR algorithm in the following lemma. Even though we only consider $d\geq 3$ in this subsection, the costs of the SOAR algorithm for $d\leq 2$ computed by Lemma 8 will be useful afterward.
@@
Suppose that $m=n$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. An algorithm ${\mathcal{A}}$ exists such that
$\mathrm{cost}_{{\mathcal{A}}}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
@@
Let ${\mathcal{A}}$ be the algorithm as per Lemma 5, and we estimate the cost of ${\mathcal{A}}$ for $d=1$, $d=2$, and $d\geq 3$ separately. Recall that an upper bound for $\mathrm{OPT}(n)$ is given by Lemma 2. For $d=1$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\right)=O\left(\int_{1}^{n}\frac{1}{\sqrt{x}}\mathrm{d}x\right)=O\left(\sqrt{n}\right)\,.$
@@
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\sqrt{\frac{\log t}{t}}\right)=O\left(\sqrt{\log n}\cdot\int_{1}^{n}\frac{1}{\sqrt{x}}\mathrm{d}x\right)=O\left(\sqrt{n\log n}\right)\,.$
@@
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{t^{1/d}}\right)=O\left(\int_{1}^{n}\frac{1}{x^{1/d}}\mathrm{d}x\right)=O\left(n^{1-\frac{1}{d}}\right)\,.$
@@
Finally, combining Lemma 8 with the lower bounds for $\mathrm{OPT}(n)$ given by Lemma 3 concludes the proof of Theorem 5.

### v3

@@ 3 Generic Framework for Leveraging Predictions of Request Locations
Given an algorithm $\mathcal{A}$} for online metric matching, there exists an algorithm $\mathcal{A}$$^{\prime}$, which is provided with a set of points $I$ before the arrival of any request, such that for all set of servers $S$ and tuple of requests $R$ with $|S|\geq|R|=|I|$, $\mathrm{cost}({\mathcal{A}}^{\prime};S,R)\leq\mathrm{OPT}(S,I)+\mathrm{cost}({\mathcal{A}};I,R)$. Moreover, if $\mathcal{A}$} is deterministic, then $\mathcal{A}$$^{\prime}$ is also deterministic.
@@
We construct the algorithm $\mathcal{A}$$^{\prime}$ as follows. Fix $S,I,R$ as required. Let $M^{S,I}$ be the matching between $S$ and $I$ with the minimum cost that matches all points in $I$. Then, we run $\mathcal{A}$} with the set of initial servers being $I$. For each new request $r$, if $\mathcal{A}$} matches $r$ to $i\in I$, then $\mathcal{A}$$^{\prime}$ matches $r$ to $M^{S,I}(i)\in S$.
@@
$\mathrm{cost}({\mathcal{A}}^{\prime};S,R)\leq\sum_{i\in I}\delta(M^{S,I}(i),i)+\mathrm{cost}({\mathcal{A}};I,R)=\mathrm{OPT}(S,I)+\mathrm{cost}({\mathcal{A}};I,R)\,$,
@@
3.1 Semi-Stochastic Setting
In this subsection, we apply Theorem 1 to the semi-stochastic setting, which enables a reduction to the stochastic setting. Specifically, we establish in the following theorem that an algorithm for the stochastic setting can be transformed into an algorithm for the semi-stochastic setting in a black-box manner. In particular, for balanced markets, this transformation preserves the competitive ratio of the algorithm up to a constant factor.
@@ 3.1 Semi-Stochastic Setting
For every distribution $\mathbb{D}$}, given an algorithm $\mathcal{A}$} for online metric matching, there exists an algorithm $\mathcal{A}$$^{\prime}$ such that $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}}^{\prime};S,n)\leq\mathrm{OPT}_{\mathbb{D}}(S,n)+\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)$ for every set of servers $S$ with $|S|\geq n$. Moreover, for balanced markets, if $\mathcal{A}$} is $\alpha$-competitive in the stochastic setting, then $\mathcal{A}$$^{\prime}$ is $(2\alpha+1)$-competitive in the semi-stochastic setting.
The rest of this subsection is devoted to proving Theorem 2. Informally, we will construct the algorithm $\mathcal{A}$$^{\prime}$ using Theorem 1, where the points in the set $I$ are independently sampled from $\mathbb{D}$. Then, we prove the guarantees of $\mathcal{A}$$^{\prime}$ by applying the triangle inequality.
We start with proving the triangle inequality for $\mathrm{OPT}(\cdot,\cdot)$.
@@
Let $M_{1}$ be the min-cost perfect matching between $S$ and $I$, and let $M_{2}$ be the min-cost perfect matching between $I$ and $R$. Note that
@@
Next, for balanced markets, we lower bound the optimal offline cost in the semi-stochastic setting in terms of that in the stochastic setting.
@@
For all distribution $\mathbb{D}$ and set of servers $S$ with $|S|=n$, $\mathrm{OPT}_{\mathbb{D}}(n)\leq 2\cdot\mathrm{OPT}_{\mathbb{D}}(S,n)$.
@@
Given a set of servers $S$ with $|S|=n$, by Lemma 5, $\mathrm{OPT}(I,R)\leq\mathrm{OPT}(I,S)+\mathrm{OPT}(S,R)$ for all sets of points $I,R$ with $|I|=|R|=n$. Taking expectations on both sides over $I\sim\mathbb{D}^{n}$ and $R\sim\mathbb{D}^{n}$, we obtain
$\mathrm{OPT}_{\mathbb{D}}(n)\leq\mathbb{E}_{I\sim{\mathbb{D}}^{n},R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(I,S)+\mathrm{OPT}(S,R)\right]=2\cdot\mathrm{OPT}_{\mathbb{D}}(S,n)\,$,
@@
Now, we are ready to prove Theorem 2.
Proof of Theorem 2..
We construct the algorithm $\mathcal{A}$$^{\prime}$ as follows. Given as input a set of servers $S$ and a tuple of requests $R$ with $|S|\geq|R|=n$, we sample a set of points $I\sim\mathbb{D}^{n}$ before the arrival of any request, and let $\mathcal{A}$$^{\prime}$ be the algorithm as per Theorem 1, which satisfies
$\mathrm{cost}({\mathcal{A}}^{\prime};S,R)\leq\mathrm{OPT}(S,I)+\mathrm{cost}({\mathcal{A}};I,R)\,.$
@@
$\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}}^{\prime};S,n)\leq\mathbb{E}_{I\sim{\mathbb{D}}^{n},R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,I)+\mathrm{cost}({\mathcal{A}};I,R)\right]=\mathrm{OPT}_{\mathbb{D}}(S,n)+\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)\,$,
which concludes the first part in Theorem 2. Furthermore, for balanced markets, if $\mathcal{A}$} is $\alpha$-competitive in the stochastic setting, then
$\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}}^{\prime};S,n)\leq\mathrm{OPT}_{\mathbb{D}}(S,n)+\alpha\cdot\mathrm{OPT}_{\mathbb{D}}(n)\leq(2\alpha+1)\cdot\mathrm{OPT}_{\mathbb{D}}(S,n)\,$,
where the second inequality holds by Lemma 6, concluding the second part in Theorem 2. ∎
4 Semi-Stochastic Setting: Euclidean Metric with Smooth Distributions
In this section, we consider the semi-stochastic setting. Moreover, we assume the metric space to be $0,1]^{d}$ with the Euclidean distance and $\mathbb{D}$ to be $\beta$-smooth [^3] The smoothness of distributions is a mild technical assumption allowing us to lower bound $\mathrm{OPT}_{\mathbb{D}}(n)$. To see the necessity, consider $\mathbb{D}$ a non-smooth distribution supported on a singleton, in which case $\mathrm{OPT}_{\mathbb{D}}(n)=0$. for some constant $\beta\geq 1$, and we apply Theorem 2 to derive algorithms for both balanced and unbalanced markets.
In Section 4.1, for constant $d\geq 1$, we give an $O(1)$-competitive algorithm for balanced markets wth the uniform distribution.
Next, we consider balanced markets with smooth distributions in Section 4.2. Recall that for general distributions, [GGPW19] give an $O(1)$-competitive algorithm for balanced markets and $d=1$, which is $O((\log\log\log n)^{2})$-competitive for general metrics. However, for $d=2$ and smooth distributions, we are only able to estimate the optimal offline cost up to a $\Theta(\sqrt{\log n})$ factor, which prevents our analysis from achieving an $o(\sqrt{\log n})$ competitive ratio. Hence, we only focus on the case with $d\geq 3$, where we present an $O(1)$-competitive algorithm.
Finally, in Section 4.3, we extend our results for balanced markets to unbalanced markets that possess at most a constant factor of excess servers, i.e., $n\leq m\leq O(n)$. We give an algorithm with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
4.1 Prelude: Balanced Markets with Uniform Distribution
In this subsection, we consider balanced markets and assume $\mathbb{D}$ to be the uniform distribution. In the following theorem, we combine Theorem 2 with the algorithm of [Kan22] to give an $O(1)$-competitive algorithm for constant $d\geq 1$.
@@ 4.1 Prelude: Balanced Markets with Uniform Distribution
Suppose that $m=n$ and $\mathbb{D}$} is the uniform distribution over $0,1]^{d}$. Then, for $d\geq 1$, there exists an $O(\sqrt{d})$-competitive algorithm in the semi-stochastic setting.
@@
We first recall the guarantees of the algorithm proposed by [Kan22].
Theorem 4 ([Kan22]).
Suppose that $m=n$ and $\mathbb{D}$} is the uniform distribution over $0,1]^{d}$. Then, for $d\geq 1$, there exists an $O(\sqrt{d})$-competitive algorithm in the stochastic setting.
Combining Theorems 2 and 4 concludes the proof. ∎
@@ 4.2 Balanced Markets
In this subsection, we assume $d\geq 3$ and consider balanced markets with smooth distributions. We prove that the SOAR algorithm is $O(1)$-competitive for $d\geq 3$ in the stochastic setting. Combining with Theorem 2 leads to an $O(1)$-competitive algorithm in the semi-stochastic setting with $d\geq 3$.
@@
Suppose that $m=n$ and $\mathbb{D}$ is a $\beta$-smooth distribution over $0,1]^{d}$ for $\beta\geq 1$. Then, for $d\geq 3$, there exists an $O(\beta^{\frac{1}{d}})$-competitive algorithm in the semi-stochastic setting.
The rest of this subsection is devoted to proving Theorem 5. By Theorem 2, it suffices to show that the SOAR algorithm is $O(\beta^{\frac{1}{d}})$-competitive for $d\geq 3$ in the stochastic setting. We compute the expected cost of the SOAR algorithm in the following lemma.
@@
Suppose that $m=n$ and $\mathbb{D}$ is an arbitrary distribution over $0,1]^{d}$. Then, there exists an algorithm $\mathcal{A}$} such that
$\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
@@
Let $\mathcal{A}$} be the algorithm as per Lemma 4, and we estimate the cost of $\mathcal{A}$} for $d=1$, $d=2$, and $d\geq 3$ separately. Recall that an upper bound for $\mathrm{OPT}_{\mathbb{D}}(n)$ is given by Lemma 1. For $d=1$,
$\mathrm{cost}_{\mathbb{D}}({\mathcal{A}};n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}_{\mathbb{D}}(t)=O\left(\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\right)=O\left(\int_{1}^{n}\frac{1}{\sqrt{x}}\mathrm{d}x\right)=O\left(\sqrt{n}\right)\,.$
@@
$\mathrm{cost}_{\mathbb{D}}({\mathcal{A}};n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}_{\mathbb{D}}(t)=O\left(\sum_{t=1}^{n}\sqrt{\frac{\log t}{t}}\right)=O\left(\sqrt{\log n}\cdot\int_{1}^{n}\frac{1}{\sqrt{x}}\mathrm{d}x\right)=O\left(\sqrt{n\log n}\right)\,.$
@@
$\mathrm{cost}_{\mathbb{D}}({\mathcal{A}};n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}_{\mathbb{D}}(t)=O\left(\sum_{t=1}^{n}\frac{1}{t^{1/d}}\right)=O\left(\int_{1}^{n}\frac{1}{x^{1/d}}\mathrm{d}x\right)=O\left(n^{1-\frac{1}{d}}\right)\,.$
@@
Finally, combining Lemma 7 with the lower bounds for $\mathrm{OPT}_{\mathbb{D}}(n)$ given by Lemma 2 concludes the proof of Theorem 5.